Write the slope-intercept equation for the line that passes through (5, -12) and (8, 12). Please show all of your work.
Answer by mananth(16949)

x1 y1 x2 y2
5 -12 8 12

x1 y1 x2 y2
5 -12 8 12

slope m = (y2-y1)/(x2-x1)
( 12 - -12 )/( 8 - 5 )
( 24 / 3 )
m= 8

Plug value of the slope and point ( 5 , -12 ) in
Y = m x + b
-12.00 = 40 + b
b= -12 - 40
b= -52
So the equation will be
Y = 8 x -52
m.ananth@hotmail.ca